OpenStudy (anonymous):

Given sin30=1/2, use trig identies to find the exact value of a. cos60 b. cos^230 c. csc pie/6 d. sec pie/3

OpenStudy (anonymous):

it has nothing to do with what sine is \[\cos(60^\circ)=\frac{1}{2}\]

OpenStudy (anonymous):

look at the unit circle on the last page of the attached cheat sheet from that you can evaluate all the above trig functions

OpenStudy (anonymous):

All of the trig ratios are inextricably linked. That's how the trig identities were developed. So, yes, cos 60 can be related to sin 30. And a cheat sheet doesn't help anyone understand the relationships between the trig ratios. The identity required in this case is\[\cos \theta = \sin \left( 90- \theta \right)\]

OpenStudy (anonymous):

How would I find \[\cos^230\]

OpenStudy (anonymous):

\[\cos ^2 30 = \cos30 \times \cos 30\]So if you apply the above identity twice you should be able to get the answer.